OpenStudy (anonymous):

word problem need help please 1. the punic wars began in 264 b.c. and ended in 146 b.b how long did the punic wars last ? 2. Roman Civil began in 509 b.c and ended in 476 a.d how long did Roman Civ last ? 3. In buffo new york the temp was 14 f in the morning. if the temp dropped 7 f what is the temp now?

OpenStudy (anonymous):

can somebody help mee

OpenStudy (anonymous):

can u help helper

OpenStudy (anonymous):

??

OpenStudy (anonymous):

one moment

OpenStudy (anonymous):

1. correction 146 bb --> 146 bc Years = 246 - 146 = 100 years 2. 509 bc to 476 ad 509 + 0 = 509 and 0 + 476 = 476 Therefore 509 + 476 = 985 3. 14 - 7 = 7 f

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

thanks one more ?

OpenStudy (anonymous):

maggie owes the candy store $35. each of 5 friends will help her pay off her debt. how much will each friend pay?

OpenStudy (anonymous):

and thats all ? for today

OpenStudy (anonymous):

35/5 = 7

OpenStudy (anonymous):

kk thanks so much

OpenStudy (anonymous):

yu are most welcome

OpenStudy (anonymous):

I have harder ones tomorrow

OpenStudy (anonymous):

shold i do harder ones today ??

OpenStudy (anonymous):

You should make sure you understand how helping did the problems above. Make sure you can do them on your own,etc. Then move on to harder ones.

OpenStudy (anonymous):

ok good idea

OpenStudy (anonymous):

r u going to be on tomorrow

OpenStudy (anonymous):

at 10 - 12 ??

OpenStudy (anonymous):

10 - 12 = -2 -- helpingtutors@gmail.com